• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 00:06:46 on January 7,2003

    Entry number 90651

    keyword=Shift Summary

    16:00 Work continues on the beam
    16:30 Access to the Hall by Arne to work on OTR while work continues in Injector
    17:15 Target Temperature Tests Underway
    18:00 OTR work finished, Vince is finishing Target Temperature Study
    18:50 15uA Beam to Hall A
    19:00 OTR Check-Out / Looks Good
    19:05 MCC Now Working On Centering The Beam On Beam Position
    21:45 Found the beam offsets are off. We need to get the new survey number!
    23:00 Final Beam Settings without Raster Setup.
    The spot with Raster is not pretty, the rastered beam is focused
    in y, so we would have to drive the raster in an eliptical shape in order
    to make a circle on the BeO target.
    23:01 An access into the Hall to fix the Hall A DAQ. (Bob Michaels)
    23:55 The ARM has found the radiation levels in the Hall to be high and
    has called in RadCon.